OpenStudy (anonymous):

Are 2 times x plus 4 all over 3 times x plus 6 and 2 over 3 equivalent expressions? Using complete sentences explain why or why not

OpenStudy (anonymous):

what is 2 times x? how do you write that?

myininaya (myininaya):

I see only one expression here? The wording seems off though.

OpenStudy (anonymous):

\[\frac{ 2x+4 }{ 3x + 6 } = \frac{ 2(x+2) }{ 3(x+ 2) } = \frac{ 2 }{ 3 }\]

myininaya (myininaya):

hmmm...I read it as \[\frac{2x+4}{3x}+6\frac{2}{3}\]

OpenStudy (anonymous):

the problem would be 2x+4/3x+6 and 2/3

OpenStudy (anonymous):

you'll want to use brackets next time ^_^. what you wrote the second time is technically: \[2x + \frac{ 4 }{ 3x } + 6\] should be (2x + 4)/(3x + 6)

OpenStudy (anonymous):

oh dang -_-